Mathematics 242 – Principles of Analysis Definitions Quiz 1 – February 1, 2013

Your Name:

A) (5) What is the definition of the union of two sets: A ∪ B?

B) (5) Let f : {A, B, C, D, E} → {1, 2, 3, 4} be a function. Is it possible for f to be one-to-one? Why or why not?